OpenStudy (anonymous):

Solve the quadratic equation. 8x2 + 16x + 8 = 0

OpenStudy (anonymous):

8x2 + 8x +8x + 8 = 0 8x(x+1) +8(x+1) = 0 (8x + 8 )(x+1) = 0 can you find the solution now????

OpenStudy (anonymous):

Solve the equation. x2 - x - 12 = 0 A) x = 4 B) x = -3 C) x = -3 or x = 4 D) x = -4 or x = 3

OpenStudy (anonymous):

x2 -4x +3x -12 = 0 x(x-4) +3(x-4) = 0 (x+3)(x-4) = 0 can you find the correct choice???